A 13.0 nC charge is at x = 0cm and a -1.4 nC charge is at x = 3 cm . At what point or points on the x-axis is the electric potential zero?
Answer #1

Electric Potential = Kq / d

due to 13.0nc + due to -1.4 nC = 0

it will be zero when at point inbetween x - 0 to 3 cm and after 5cm

so taking Case1:

k * 13.0nC / d   +    k x -1.4nC / (d-3 ) = 0

- 39 + 13.0d = 1.4d

d = 3.36 cm .........Ans

And Case2:

k * 13.0nC / d   +    k x -1.4nC / (3-d ) = 0

39 - 13.0d = 1.4d

d = 2.71 cm ......Ans

x = 2.71 cm , 3.36 cm .....final answer

I hope help you !!